# What computer science has to say about the simulation hypothesis

### v4

@@ 2.5 The Reverse Physical Church-Turing Thesis
Loosely speaking, the PCT says that the dynamics of any universe that we are considering can be computed on a UTM. One can “reverse” the requirement that a universe obey the PCT, to produce the requirement that a universe contains a UTM in it. If it obeys such a reversed PCT, a universe could implement all TMs.
One might imagine that the reversed PCT could be formalized in a way that among other things requires that a universe’s computer $N$ evolves independently of the rest of that universe. However, in general that is not possible — we need to allow the beings running the computer to provide it information in order to characterize its computational power, which means that that computer $N$ does not evolve autonomously as a UTM. So we cannot impose this simple version of a reverse PCT. On the other hand, we can require that $N$ effectively implements a UTM. This is done as follows:
@@
such that
  1. 1.
    those three functions are well-defined for all pairs of TM index $k$ and $y\in{\mathbb{B}}^{*}$ such that $T^{k}$ halts on input string $y$;
  2. 2.
    for all such $(k,y)$,
@@
I say that three functions $\widehat{{\mathcal{N}}},\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}}$ all taking $\mathbb{N}$$\times{\mathbb{N}}\rightarrow{\mathbb{N}}$ have the RPCT properties for $g$ if they obey the properties listed in Definition 4.
Note that there is no concern in  Definition 4 for the form of the functions $\widehat{{\mathcal{N}}},\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}}$ for any argument $(k,y)$ where $T^{k}$ does not halt on input string $y$. They may even be undefined for such an argument. Moreover, a priori, the three functions in Definition 4 could be uncomputable for some $(k,y)$ pair even when $T^{k}$ does halt on input string $y$. Since no assumption is made in this paper that any universe / alien beings / computer actually constructs those functions, that wouldn’t matter.
Nonetheless, it is an important special case when those three functions are partial computable, in a specific way. Abusing terminology, I say that the computable RPCT holds for $V$ iff all three functions $\widehat{{\mathcal{T}}}(k,y),\widehat{{\mathcal{W}}}(k,y),\widehat{{\mathcal{N}}}(k,y)$ are computable over the set of all pairs of arguments $(k,y)$ such that $T^{k}(y)$ halts. For simplicity, unless specified otherwise, throughout this paper I will assume that whenever the RPCT holds, in fact the computable RPCT holds.
Broadly speaking, the RPCT says that the system $N$ operates like a UTM for all pairs $(k,y)$ such that $T^{k}(y)$ is defined, where $k$ and/or $y$ may be encoded in some degrees of freedom in $w_{0}$ rather than directly in $n_{0}$. That freedom to have $k$ and / or $y$ specified in the environment of the computer $N$ allows that computer to do things like observe its environment to retrieve the input string for a computation from its environment before running that computation. (Note that this freedom is just like the property of free simulation, defined above.)
Exploiting this freedom, I say that the special case of pristine RPCT holds if for all $k,y$, the RPCT holds with
@@
Eqs. 6 and 7 mean that the physical system implementing the UTM is initialized with the precise TM that it is to supposed to implement, but not the actual data that TM will be run on. In general, that data (the value of $y$) would be transferred into the computer $N$ at a subsequent iteration, after this initialization of the UTM, but before the UTM starts to run. As an example, this is what would occur if we were to initialize a laptop with a universe-simulating program, with the precise data that program is to run on fed into the laptop before starting the simulation program. (See Section 2.6.) In this paper, when assuming the RPCT I will not also implicitly assume that the pristine RPCT holds; it only holds if I explicitly say so.
@@
Despite the popularity of the RPCT among researchers, even if the PCT holds in our specific universe, that does not mean that the RPCT has to hold as well. In particular, cosmological considerations could prevent it from holding [43]. So the results below specific to universes that obey the PCT and / or universes that obey the RPCT might not apply to our specific universe. However, approximations of the analysis might apply even if the PCT and/or RPCT don’t hold exactly in our universe, depending on how precisely they are violated.
Even without worrying about the laws of physics in our universe, one might suppose that the RPCT is logically impossible, and so cannot hold in any universe. After all, the RPCT requires that the set $X$ of physical variables of a universe decomposes as $X=W\times N$ where $N$ is a physical system that implements a UTM. In other words, a proper subset of the spatial degrees of freedom of the universe constitutes a physical structure $N$ capable of implementing any TM. But the PCT thesis supposes that such an $N$ would be able to simulate the dynamics of all of $X$. So this $N$ would have to be able to simulate itself, at the same time as it is also simulating the entire rest of $X$, outside of $N$. This would seem to imply a contradiction, that $N$ is more computationally powerful than $N$ is.
@@
One might be suspicious of this argument though, since it is quite similar to the arguments that were made in the last century that no physical system can make an extra copy of itself without destroying itself. (These arguments that copying required destruction of the original were used to make the case that the common definition of life involving replication must be wrong, or at least deficient.) Responding to these arguments, Von Neumann designed his “universal constructor” in a cellular automaton setting. This demonstrated explicitly that it is possible for a system to copy itself without harming itself in the process. [^12] It is interesting to note that Von Neumann’s proof of this property of his universal constructor is essentially identical to Kleene’s earlier proof of his second recursion theorem — and that it is hard to imagine that Von Neumann did not know of Kleene’s earlier result, despite not citing it in his work.
2.6 Sketch of how Newtonian universes might obey the PCT and RPCT
This section discusses how both the RPCT and PCT could be obeyed in our physical universe. I first mention the literature on this issue. Most of that literature presumes that the physical experimentalist can access real value variables to infinite precision. So I present a new example of a system that obeys both the PCT and the RPCT, which only involves Newtonian mechanics.
There is a very large literature showing how to design Newtonian physical systems that are computationally universal [48, 16, 58]. Those establish RPCT, and since Newtonian mechanics is computable, also the PCT, since we’re treating the entire universe, outside and inside of $Y$, as Newtonian.
Typically such examples from the literature rely on the ability of the person using a physical system as a computer to manipulate and observe real-valued physical variables in that system to arbitrarily fine precision. One might object that this is beyond the capabilities of humanity, both now and in the foreseeable future. However, it is straight-forward to design a Newtonian system that is computationally universal but doesn’t require the ability to access physical variables to arbitrary resolution. The rest of this subsection demonstrates one such design.
Fix some finite set of point masses evolving through a perfect vacuum under mutual gravitational attraction. Suppose there no other forces coupling them after the initial time $t=0$. So in particular, after $t=0$ there are no interactions between those point masses and the universe external to them. Let $\Gamma$ be the phase space of that set of point masses.
Identify the initial IDs of a UTM at time 0 with the set $\mathcal{S}(0)$ of points in a countably infinite grid over $\Gamma$. Assume as well that at all of those grid points, at least one of the point masses has nonzero momentum. Write $\gamma_{y}(t)$ for the phase space position at time $t>0$ of the system if it was at grid point $y\in\mathcal{S}(0)$ at time 0. Write the set of all phase space positions given by evolving the positions in $\mathcal{S}(0)$ forward from time 0 to time $t$ as $\mathcal{S}(t)$.
Define the infinite set of discrete times $\{t_{i}\}={\mathbb{Z}}^{+}$. Also, choose the grid points $\mathcal{S}(0)$ so that for any two distinct time-0 grid points, $y,y^{\prime}\neq y\in\mathcal{S}(0)$, the associated curves they trace out as they evolve through $\Gamma$, $\{\gamma_{y}(t):t>0\}$, and $\{\gamma_{y^{\prime}}(t):t>0\}$, never cross one another. (For example, one could do this by having different momentum vectors of the centers of mass of all the points in $\mathcal{S}(0)$.) This ensures that there is no overlap between any two sets $\mathcal{S}(t_{i})$ and $\mathcal{S}(t_{j})$ for $i\neq j$.
Now simply identify each of the initial grid points $y\in\mathcal{S}(0)$ with a different initial ID of some fixed UTM, and the subsequent points it goes through at the times $t_{i}$, $\gamma_{y}(t_{i})$, with the subsequent IDs of the UTM that started with the ID $y$. Since $\mathcal{S}(t_{i})\cap\mathcal{S}(t_{j})=\varnothing$ for $i\neq j$, we are guaranteed that there is no ambiguity of what ID should be identified with a given position in $\Gamma$.
Despite this though, in general there will be some time $t_{i}>0$ after which there is a growing set of grid points that are assigned the same ID as some other grid point. (This is due to merging of trajectories through the set of all IDs of the UTM.) There will also be times after which there will some IDs that do not correspond to any of the grid points. (This is due to the fact that some IDs of the UTM cannot arise after a certain number of iterations, no matter what input tape the UTM starts with.) Moreover, in general the same ID will be assigned to multiple points in $\Gamma$. Finally, note that if $\gamma_{y}(t_{i})$ is assigned to an ID which is a halt state, then for all $j>i$, $\gamma_{y}(t_{j})$ is assigned that same halting ID, $\gamma_{y}(t_{i})$.
In any case, this system obeys the RPCT by construction, if we identify $N$ with the UTM whose dynamics is given by the trajectories $\{\gamma_{y}(t_{i}):y\in\mathcal{S}(0),t_{i}\in{\mathbb{Z}}^{+}\}$. More formally, for all $k,y$ as defined in Definition 4, write $\langle k,y\rangle$ for the initial ID of our UTM (i.e., the initial contents of its input tape) that will cause it to compute $T^{k}(y)$ and then halt. It suffices to establish the RPCT by choosing
  • $\widehat{{\mathcal{T}}}(k,y)$ is the halting time for our UTM to compute $T^{k}(y)$ when initialized to the ID $\langle k,y\rangle$;
  • $\widehat{{\mathcal{W}}}(k,y)$ is arbitrary;
  • $\widehat{{\mathcal{N}}}(k,y)$ is $\langle k,y\rangle$.
This establishes the RPCT, as claimed. Moreover, since Newtonian mechanics is computable and we’re treating the entire universe, both inside and outside of $\Gamma$, as Newtonian, the PCT holds automatically, as claimed.
See Appendix D for discussion of some subtleties of this scheme for implementing a universe $V$ that obeys the full CT using Newtonian mechanics.
@@ 3.2 Why the simulation lemma might seem to preclude self-simulation
The simulation lemma tells us that if there are some super-sophisticated aliens in a universe $V$ in which the RPCT holds, and if our universe obeys the PCT, then it’s possible that we are simulations in a computer that the aliens are running. (Of course, we would never know it if they were.) On the other hand, suppose that as many have argued the PCT does not hold in our universe, because its evolution cannot be evaluated on a TM. In such a case, even if the RPCT holds in the universe that the aliens inhabit, there is no guarantee that we may be in a simulation that they are running. This formalizes the intuitive idea that our universe has to be “sufficiently simple, computationally speaking” in order for us to be simulations in a computer of some super-sophisticated aliens.
Note though that even if the conditions in the simulation lemma hold, that lemma doesn’t in some sense specify what argument $(\Delta t^{\prime},w^{\prime}_{0},n^{\prime}_{0})$ to simulate. In particular, it does not tell us what value $n^{\prime}_{0}$ the computer $N^{\prime}$ that $N$ is simulating would need to start with in order for that simulation to give the dynamics of $V$ itself. To understand the implications of this, suppose that: i) the RPCT holds for our universe, so we have a computer $N$ we can run to simulate the evolution of any universe that obeys the PCT; ii) our universe itself obeys the PCT. One might suppose that when these two conditions are met, Lemma 1 implies that we can run a simulation of our own universe, including ourselves. However, the simulation lemma provides no means for us to initialize our computer $N$ in order to run a simulation of ourselves. It provides no way to have a universe $V$ determine how to simulate itself.
@@
We have no guarantee that the solution $n^{\prime}$ to this fixed point equation is computable. In fact, a priori, one might suspect that it is possible for there not to be any solution to Eq. 12 whatsoever, computable or otherwise. In light of Lemma 1, if that were the case, it would mean that the conditions for the simulation lemma cannot be met, i.e., that it is not possible for a universe $V$ to obey both the PCT and the RPCT.
@@
Finally, none of the analysis above establishes that it is even logically possible for a universe to obey both the PCT and the RPCT. It may be mathematically impossible to meet the conditions for Lemma 2.
@@ 4 The self-simulation lemma
In this section I begin by proving that despite the arguments in Section 3.2, in fact self-simulation is possible. The key is to have a time delay between the future moment for which the state of the universe is being simulated on the one hand, and on the other hand, the moment when the computer’s simulation of that future state completes.

### v5

@@ 2.5 The Reverse Physical Church-Turing Thesis
Loosely speaking, the PCT says that the dynamics of any universe that we are considering can be computed on a UTM. One can “reverse” the requirement that a universe obey the PCT, which results in the requirement that a universe contains a UTM in it. If it obeys such a reversed PCT, a universe could implement all TMs. (Note that this does not mean that the universe itself is a universal TM; it could have super-Turing power, in which case it is not any type of TM at all.)
One might imagine that the reversed PCT could be formalized in a way that among other things requires that a universe’s computer $N$ evolves independently of the rest of that universe. However, in general that cannot be assumed in the analysis below — we will need to allow the beings running the computer to provide information to that computer after it starts running. (As an example, this is the case whenever one runs a computer program that gets input from the user or an external disk drive after it starts running.) This means that that computer $N$ does not evolve autonomously as a UTM. So we cannot impose this simple version of reverse PCT.
On the other hand, we can require that $N$ effectively implements a UTM. This is done as follows:
@@
that are well-defined for all TM indices $k$ and all $y\in{\mathbb{B}}^{*}$ such that $T^{k}$ halts on input string $y$, where for all such $(k,y)$,
@@
I say that three functions $\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}},\widehat{{\mathcal{N}}}$ all taking $\mathbb{N}$$\times{\mathbb{N}}\rightarrow{\mathbb{N}}$ have the RPCT properties for $g$ if they obey the properties listed in Definition 4. Abusing terminology, I also say that the computable RPCT holds for $V$ iff all three functions $\widehat{{\mathcal{T}}}(k,y),\widehat{{\mathcal{W}}}(k,y),\widehat{{\mathcal{N}}}(k,y)$ are computable for all pairs of arguments $(k,y)$ such that $T^{k}(y)$ halts. (They may be undefined for other $(k,y)$, i.e., they may be only partially computable.) Unless specified otherwise, throughout this paper I will assume that whenever the RPCT holds, it’s computable.
Broadly speaking, the RPCT says that the system $N$ operates like a UTM for all pairs $(k,y)$ such that $T^{k}(y)$ is defined, where $k$ and / or $y$ may be encoded in some degrees of freedom in $w_{0}$ rather than directly in $n_{0}$. (That freedom to have $k$ and / or $y$ specified in the environment of the computer $N$ allows that computer to do things like observe its environment to retrieve the input string for a computation from its environment before running that computation.)
I say that the pristine RPCT holds if for all $k,y$, the RPCT holds with
@@
Eqs. 6 and 7 mean that the physical system implementing the UTM is initialized with the precise program that it is to supposed to implement, but not the actual data that program will be run on. In general, that data (the value of $y$) would be transferred into the computer $N$ at a subsequent iteration, after this initialization of the UTM, but before the UTM starts to run. As an example, this is what would occur if we were to initialize a laptop with a universe-simulating program, with the precise data that program is to run on is fed into the program before the laptop starts evaluating that simulation. (See Section 2.6.)
Note that in the definition of the pristine RPCT I don’t introduce the extra notation that would be needed if the extra information concerning $y$ that $N$ would use is contained in a proper subset of the variables in $W$. Instead I simply equate the value $y$ with $w\in W$ in its entirety. This is just for pedagogical simplicity. In any case, when assuming the RPCT below I will not implicitly assume that the pristine RPCT holds unless I explicitly say so.
@@
Despite the popularity of the RPCT among researchers, even if the PCT holds in our specific universe, that does not mean that the RPCT has to hold as well. In particular, cosmological considerations could prevent it from holding [38]. So the results below specific to universes that obey the PCT and / or universes that obey the RPCT might not apply to our specific universe. However, approximations of the analysis might apply even if the PCT and/or RPCT don’t hold exactly in our universe, depending on how precisely those properties are violated.
Even without worrying about the laws of physics in our universe, one might suppose that the RPCT is logically impossible in our cosmological universe, if the PCT holds for our universe. After all, the RPCT requires that the set $X$ of physical variables of a universe decomposes as $X=W\times N$ where $N$ is a physical system that (once it reads in the data $y$) implements a UTM. In other words, a proper subset of the spatial degrees of freedom of the universe constitutes a physical structure $N$ capable of implementing any TM. But the PCT thesis supposes that such an $N$ would be able to simulate the dynamics of all of $X$. So this $N$ would have to be able to simulate itself, at the same time as it is also simulating the entire rest of $X$, outside of $N$. This would seem to imply a contradiction, that $N$ is more computationally powerful than $N$ is.
@@
One might be suspicious of this argument though, since it is quite similar to the arguments that were made in the last century that no physical system can make an extra copy of itself without destroying itself. (These arguments that copying required destruction of the original were used to make the case that the common definition of life involving replication must be wrong, or at least deficient.) Responding to these arguments, Von Neumann designed his “universal constructor” in a cellular automaton setting. This demonstrated explicitly that it is possible for a system to copy itself without harming itself in the process. [^8] It is interesting to note that Von Neumann’s proof of this property of his universal constructor is essentially identical to Kleene’s earlier proof of his second recursion theorem — and that it is hard to imagine that Von Neumann did not know of Kleene’s earlier result, despite not citing it in his work. In modern parlance, computer viruses are mathematically possible – which is perhaps comforting, since in fact they exist.
2.6 Sketch of how a portion of our cosmological universe could obey the PCT and RPCT
In the rest of this subsection I sketch how the framework defined above might apply to a portion of our actual cosmological universe, evolving according to the laws of physics as we currently understand them.
Choose some particular inertial frame to describe physical systems in our universe. (Sticking to this specific inertial frame will allow us to circumvent the complications of special relativity.) Let $Y$ be a special subset of the space-time of our cosmological universe defined by a cylinder in our chosen inertial frame, a cylinder whose base consists of a $3$-dimensional region, $Y_{0}$. $Y_{0}$ contains an infinite number of accessible degrees of freedom, and all of the points in $Y_{0}$ have the same value of time in our chosen inertial frame, $t_{0}$, a time which is long after the Big Bang. (Having $t_{0}$ be long after the Big Bang will allow us to circumvent the complications of cosmological issues involving general relativity and the boundary conditions of our cosmological universe, and will allow us to talk about initializing various physical variables in a computer that resides in $Y$.) For completeness, suppose that $Y$ extends infinitely to the future of time $t_{0}$. Discretize time (in our co-moving inertial frame) going forward from $t_{0}$, writing the time at the end of each interval as $t_{i}$, $i\in{\mathbb{N}}$, and the associated state of our region as $Y_{i}:=Y_{t_{i}}$.
Let $V=W\times N$ be a coarse-graining of that $3$-dimensional region defining $Y$ into a set of classical (non-quantum) bins, where $W$ is finite and N{N} is countably infinite. As described above, it is useful to parameterize N{N} as the set $\mathbb{B}$$^{*}$ of all finite strings of bits. Suppose that the variables given by that coarse-graining follow deterministic evolution. This requires in particular that $V$ be physically isolated from the rest of the universe, and so they collectively obey Hamilton’s equations. Note that by considering the dynamics of $V$, we can circumvent the complications of quantum mechanics, and in particular of extending the PCT thesis to the quantum realm.
Since Hamilton’s equations are computable, and $X_{t}$ obeys Hamilton’s equations, the PCT is obeyed by the dynamics of $X_{t}$. We also assume the projected dynamics down to $N$ implements a UTM (though perhaps it does so slowly, limited by the speed of light among other factors). Therefore the RPCT is also obeyed by the dynamics of this universe, $V$.
Given this general setup, choose $w_{0}\in W$ arbitrarily. Also pick an initial ID $n_{0}\in N$ of the UTM that has an arbitrary (blank-delimited) finite string on the input tape, with all other tapes (if any) initially fully blank, and the pointers of the UTM in their initialized positions.
Choose the dynamics of the entire universe $V$ so that the first thing that happens when it starts evolving from its initial state is that a copy of $w_{0}$ is placed after the string $n_{0}$ on the input tape of the (physical system implementing) the UTM. Physically, this could reflect the process in which the being who is going to run the simulation program on the computer feeds that computer the input data it needs concerning the external universe. Alternatively, it could reflect the computer being coupled to an observation device which observes that data and provides it to the computer. After that the dynamical laws (Hamilton’s equations in this universe) evolve both $W$ and $N$, perhaps via a coupled dynamics, or perhaps completely independently of one another.
As an aside, it is unusual in standard physics to consider state spaces $\mathbb{B}$$^{*}$ (which is the space we assign to $N$). However, such spaces arise naturally when modeling a discrete computer, e.g., a laptop, whose memory is always finite — but can be expanded by an arbitrary finite amount an arbitrary number of times. Viewing that memory as $M$, we can formalize its dynamic expansion in terms of dynamics over the uncountably infinite space $\mathbb{B}$$^{\infty}$, by “embedding” $N$ into $\mathbb{B}$$^{\infty}$ as the indexed set of spaces $N_{t}={\mathbb{B}}^{K(t)}$ for some function $K:{\mathbb{Z}}^{+}\rightarrow{\mathbb{N}}$. We define this function $K(t)$ recursively. First, specify $k(0)$ to be some arbitrary finite counting number. Then if at time $t$ there are any bits in ${\mathbb{B}}^{\infty}\setminus N_{t}$ whose state at $t+1$ is causally dependent on the state of $n_{t}\in N_{t}$, $K(t)$ increases enough to include those bits. (Note the implicit assumption that at every iteration $t$, there are only a finite number of such bits whose state may depend on $n_{t}$.)
See Appendix D for discussion of some subtleties of the scheme outlined above for implementing a universe $V$ in our actual, cosmological universe.
@@ 3.2 Why the simulation lemma might seem to preclude self-simulation
The simulation lemma tells us that if there are some super-sophisticated aliens in a universe $V$ in which the RPCT holds, and if our universe obeys the PCT, then it’s possible that we are simulations in a computer that the aliens are running. (Of course, we would never know it if that were the case.) On the other hand, suppose that, as many have argued, the PCT does not hold in our universe, because our universe’s evolution cannot be evaluated on a TM. In such a case, even if the RPCT holds in the universe that the aliens inhabit, there is no guarantee that we may be in a simulation that they are running. This formalizes the intuitive idea that our universe has to be “sufficiently simple, computationally speaking” in order for us to be simulations in a computer of some super-sophisticated aliens.
The situation becomes more philosophically interesting if we consider the case that the universe being simulated, $V^{\prime}$, obeys the PCT, and in addition is $V$ itself. This would mean in particular that $N$ is being used to simulate the evolution of $N$.
To investigate this possibility, note that even if the conditions in the simulation lemma hold, that lemma doesn’t in some sense specify what argument $(\Delta t^{\prime},w^{\prime}_{0},n^{\prime}_{0})$ to simulate. In particular, it does not tell us what value $n^{\prime}_{0}$ the computer $N^{\prime}$ that $N$ is simulating would need to start with in order for that simulation to give the dynamics of $V$ itself. To understand the implications of this, suppose that: i) the RPCT holds for our universe, so we have a computer $N$ we can run to simulate the evolution of any universe that obeys the PCT; ii) our universe itself obeys the PCT. Lemma 1 would seem to imply that when those two conditions are met, we can run a simulation of our own universe, including ourselves. The issue is that given our definition of “simulate”, we need to specify the precise initial state of the universe being simulated. The simulation lemma provides no means for us to choose that initial state for $N$ to run a simulation of $V$, nor to initialize our computer $N$ in order to run a simulation of ourselves. It provides no way to have a universe $V$ determine how to simulate itself.
@@
We have no guarantee that the solution $n^{\prime}$ to this fixed point equation is computable. In fact, a priori, one might suspect that it is possible for there not to be any solution to Eq. 12 whatsoever, computable or otherwise. In light of Lemma 1, if that were the case, it would mean that the conditions for the simulation lemma cannot be met, i.e., that it is not possible for a universe $V$ to obey both the PCT and the RPCT. Indeed, none of the analysis above establishes that it is logically possible for a universe to obey both the PCT and the RPCT. That means that it may be mathematically impossible to meet the conditions for Lemma 2.
@@ 4 The self-simulation lemma
In this section I begin by proving that despite the arguments in Section 3.2, in fact self-simulation is possible. The key is to have a time delay between the future moment for which the state of the universe is being simulated on the one hand, and when the computer simulation of that future state completes on the other hand.